OpenStudy (anonymous):

a deck of cards has 4 purple , 4 navy, and 4 violet cards. you pick four cards from the deck. cards a returned to the deck after they are picked. p(the first card is navy, the second card is navy,the third card is purple, and the fourth card is violet)

OpenStudy (anonymous):

P(navy) = 1/3 P(purple) = 1/3 P(violet) = 1/3 since every draw is independent with each other (because the cards are returned), then you just multiply the probs. P(navy) * P(navy) * P(purple) * P(violet) = (1/3) ^ 4 or 1/81